# Nocturnal Eye-Inspired Liquid-to-Gas Phase Change Soft Actuator with Laser-Induced Graphene: Enhanced Environmental Light Harvesting and Photothermal Conversion

### v2

@@ Abstract
Robotic systems’ mobility is fundamentally constrained by their power sources and wiring requirements. While electrical actuation systems have achieved autonomy through battery power and wireless control, pneumatic actuators remain tethered to air supply sources. Liquid-to-gas phase change actuators utilizing low-boiling-point liquids offer a potential liquid, though they typically require substantial thermal input through heating elements that maintain electrical dependencies. External heat sources, particularly light energy, present an alternative for terrestrial applications. However, despite their optical transparency, silicone-based materials have high volumetric heat capacity and low thermal conductivity, limiting efficient photothermal energy transfer. Previous attempts to address this through graphene or metallic powder incorporation have compromised material properties, including reduced transparency and altered elastic moduli. Inspired by the tapetum lucidum structure found in nocturnal animals’ eyes for efficient light utilization in low-light conditions, this study proposes a novel anisotropic bilayer soft actuator incorporating Laser-Induced Graphene (LIG) on the inner surface of the light-irradiated silicone layer. This creates an anisotropic structure with enhanced photothermal conversion capabilities while maintaining the advantageous properties of silicone. Comparative analysis demonstrates that the proposed actuator exhibits significantly higher photo-induced bending efficiency than conventional silicone-based actuators. The response time (T63) improved by 54%, decreasing from 142 seconds for pure silicone to 65 seconds, with recovery response time showing a 48% improvement. This design maintains silicone’s transparency and flexibility while utilizing LIG, which can be fabricated under ambient conditions, facilitating manufacturing and diverse applications.
@@ 1 Introduction
Currently, most robotic actuators rely on electrical energy. For example, although pneumatic actuators use air as the driving source, they rely on electrical energy to control and supply pressed air [1][2]. Therefore, they cannot be considered completely independent of electrical energy. Throughout history, humans have utilized various forms of energy to enrich their lives. Aiming to use energy sources other than electrical energy to drive soft robots can significantly contribute to diversifying actuation methods in robotics.
Several methods have been proposed for using heating methods, such as laser irradiation [3][4], microwave [5], heatsink [6] and electromagnetic [7][8]. Kim et al. developed an electrothermal pneumatic actuator using nanofiber mats, achieving a lightweight, soft actuator that can operate without external devices by utilizing liquid evaporation [9]. However, this approach faces issues such as high power dependency and slow response times.
Various proposals have been made, starting with actuators like pouch motors ([10])([11])([12])([13])that utilize liquid-to-gas phase change, including pneumatic actuators driven by ambient temperature ([14]). However, these actuators operate relatively slowly, with many functioning at frequencies below 0.1Hz. This is due to significant losses in converting environmental energy into driving energy, heat transfer, and thermal conductivity issues. Sogabe et al.[15] developed a small soft actuator that operates in environments rich in thermal energy, such as warm water. This innovation improved the actuation time for bending to around one second using only environmental energy. However, this actuator leverages the thermal properties of the surrounding water (specific heat capacity of 4186 $\rm{J/kg\cdot K}$[16]) and is thus limited to underwater use.
For broader environmental applications.., systems capable of operating on land are required. A solar-driven actuator presents a promising approach to solving these issues [17]. Shao et al. proposed an actuator that absorbs environmental light energy through blackbody radiation by mixing graphite into silicone, converting it into thermal energy, and driving it using the liquid-to-gas phase change (methanol)[18]. However, this method results in the loss of actuator transparency due to the use of graphite. Adding powders like graphite reduces Young’s modulus, leading to constraints on the actuator’s shape and functionality.
Therefore, it is crucial to develop a mechanism that efficiently absorbs environmental energy, enabling actuation even with highly transparent silicone actuators. We introduce mechanisms inspired by biological systems that efficiently absorb environmental energy into the actuator.
@@
Nocturnal animals need to hunt and escape predators under the moon’s and stars’ faint light. To efficiently transmit this minimal light to the retinal photoreceptors, they possess a unique structure called the tapetum lucidum (Fig. 1A). This structure captures light that passes through the photoreceptors and reflects it to the retinal layer, again stimulating the photoreceptors. Inspired by the relationship between the tapetum and the retina, which ’recovers and reuses light that would otherwise escape, our study proposes the introduction of a new light-reusing bilayer structure that enables liquid-to-gas phase change, which is inefficient with silicone layers alone.
In this study, we employed Laser-Induced Graphene (LIG) as the photothermal conversion element for the "tapetum" layer. In recent years, Laser-Induced Graphene (LIG) has emerged as a prominent research focus in direct laser writing techniques, where organic materials such as polyimide are transformed into graphene-like structures. LIG is characterized by the formation of a porous three-dimensional graphene network structure, produced through rapid thermal carbonization (pyrolysis) via laser irradiation. Lin et al. proposed a method to directly graphenize polyimide film surfaces using $\rm{CO_{2}}$  laser irradiation[19]. Subsequently, Chyan et al.[20] demonstrated the versatility of LIG fabrication on diverse substrates, including textiles, paper, and food products, through repeated laser exposure, thereby accelerating its application in flexible devices. LIG has garnered significant attention as a photothermal conversion material that requires no electrical connections or external wiring. It exhibits efficient light-to-heat conversion properties under laser or lamp irradiation, demonstrating transferability to elastomeric substrates. For instance, Dallinger et al. developed an actuator by transferring LIG onto PDMS-based hydrogel sheets, where the generated heat-induced humidity changes in the sheet, resulting in strain-induced bending[21]. These characteristics have positioned LIG as a promising material for applications in bioelectronics and soft actuators[22].
Building upon these advances, we focused on expanding LIG’s applications to liquid-to-gas phase change actuators. To enable efficient light-driven actuation, we added a LIG layer to conventional silicone soft actuators[15]. We propose a novel anisotropic bilayer structure with a LIG layer on the light-incident surface inside highly transparent silicone. This design, inspired by the eyes of nocturnal animals, enables efficient light collection and photothermal conversion. The converted thermal energy is transferred to the internal fluid, causing deformation of the silicone actuator and resulting in controlled bending motion. In this study, we present the fabrication process of the proposed actuator and evaluate its performance characteristics. Furthermore, to demonstrate the practical applicability of our system, we developed a multi-finger robot where multiple units of these novel actuators operate synchronously in response to light stimuli.
@@ 2.1 Concept
Figure 1 illustrates the conceptual design of our proposed actuator. While previous phase-change actuators typically employed homogeneous structures, our novel approach focuses on the layered wall architecture for efficient light absorption. As previously discussed, biological eyes utilize the tapetum to reflect escaped light, thereby increasing photoreceptor stimulation. Similarly, our design incorporates a Laser-induced graphene (LIG) layer to capture light not absorbed by the conventional silicone elastomer layer, converting it to thermal energy through photothermal conversion to efficiently vaporize and expand the internal low-boiling-point liquid (Fig. 1A).
The actuation mechanism of the proposed device is depicted in Fig. 1B. The soft actuator deforms by vaporizing an internal low-boiling-point liquid controlled by an external light source. The actuator is designed with asymmetric silicone wall thicknesses, with the thinner wall on the graphene layer side to induce directional bending toward the LIG layer. The operating temperature was set at 33C, above room temperature (25C), and we selected an inert fluorinated liquid (Opteon SF33, Chemours Company, base compound structure: $\rm{Cis-CF_{3}CH=CHCF_{3}}$) as the working fluid.
Figure 1C shows a photograph of the fabricated actuator. The actuator has an overall length of 20 mm with a flattened elliptical cross-section measuring approximately 5 mm along the central axis and 4 mm along the minor axis. As illustrated in Fig. 1D, our proposed actuator achieves a layered structure using LIG without directly incorporating thermally conductive materials or photothermal converters into the silicone walls. This approach facilitates efficient heat transfer and maintains the transparency of the silicone elastomer, enabling broader applications of the elastomer-based system.
@@ 2.2.1 LIG Transfer Process onto Silicone
The LIG patterns are generated on Kapton tape (3 M Polyimide Film Tape 5413, 3 M Company, Minnesota, USA) adhered to a slide glass. The polyimide base of the Kapton tape enables LIG formation through laser irradiation. We utilize a laser system (xTool F1, Makeblock Co., Ltd., Guangdong, China) with specific parameters (200 mm/s speed, 40% power, unidirectional) for LIG patterning[23].
@@ 2.3 LIG-transferred soft actuator fabrication process
Figure 2 illustrates the fabrication process of a soft actuator with the LIG layer positioned adjacent to the liquid interface. The fabrication process consists primarily of five steps: (i)–(v). Initially, a LIG pattern (2.6 mm $\times$ 40 mm) is created using the aforementioned laser parameters. The pattern is then positioned and secured within the mold (detailed in Supplementary Data S1) as shown in (i-a). Subsequently, a metallic rod (2.5 mm in diameter) is inserted into the mold, and silicone is introduced and cured, as depicted in (ii). In step (iii), the tubular silicone structure with the integrated LIG layer is carefully removed from the mold. Because the LIG layer is initially present on the external surface, the tube is inverted in step (iv) to position the LIG layer on the interior. In step (v), uncured silicone is introduced into one tube end to create a seal, resulting in an actuator with one open terminus. The actuator is completed by introducing 40 µL of a low-boiling-point liquid (Opteon SF33, The Chemours Company, Delaware, USA) into the internal cavity and sealing the open end with a silicone-coated clip.
Additionally, the silicone sheets (10 mm square with transferred LIG patterns) were fabricated using the same process with molds adjusted to achieve a thickness of 1.0 mm. These molds were fabricated using a Formlabs Form 3 printer (Formlabs Inc., Somerville, MA, USA). Clear Resin v4 (Formlabs Inc.) was used as the mold production resin material.
We examined the temperature response characteristics to investigate whether placing the silicone layer on the exterior and the LIG layer on the interior would affect LIG’s photothermal conversion efficiency. A reflector lamp (RF100 V54 WD, 100 V, 54 W; Panasonic Corporation, Osaka, Japan) was used as the light source, and temperature changes on the silicone sheet were measured using a thermal camera (OPTPI64 ILTO15 T090, Optris GmbH, Berlin, Germany).
Our findings revealed that positioning LIG on the interior actually accelerated the temperature rise at the liquid-contact interface (Supplementary Fig. 1). The initial temperature change rate during the first 5 seconds after light irradiation was 0.71 K/s when irradiated from the silicone side, compared to 0.66 K/s when irradiated from the LIG side (reference value: approximately 0.13 K/s for silicone-only sheet). This phenomenon can be attributed to silicone’s high volumetric heat capacity and low thermal conductivity, which impede the efficient transfer of heat generated by LIG to the liquid-contact interface.
@@ 2.4 Mechanical and chemical durability of LIG-transferred silicone
Figure 3 illustrates the silicone sheet (10 mm $\times$ 10 mm) with a transferred LIG layer, as well as the condition of the LIG after being subjected to 20 cycles of stretching and, relaxation, and twisting, respectively. In both cases, no delamination of the LIG layer was observed.
The resistance of the actuator to the liquid was evaluated by immersing the LIG-transferred sheet into a conical tube filled with Opteon SF33. The actuator was agitated for 60 seconds. After the process, the sheet was examined, and no detachment or degradation of the LIG layer was observed, confirming its stability under these conditions (Fig. 4).
@@ 3 Heat transfer analysis
Here, $\sigma=5.67037442\times 10^{-8}~{}\mathrm{W\cdot m^{-2}\cdot K^{-4}}$ is the Stefan-Boltzmann constant; $\theta_{s}$ is the temperature of the silicone wall; $A$ is the heat transfer area; and $\epsilon_{h}$ and $\epsilon_{s}$ are the emissivities of the heat source and the silicone wall, respectively ([24]). The thermal conductivity of the silicone, $\lambda_{s}$, and its thickness, $x$, are given in Table 1 [25].
@@
The parameters used in the simulation are summarized in Table 1[26].[27]
@@
For comparison with simulation results, we measured the temperature at the liquid contact surface using a thermal camera with silicone sheets of the same thickness as in the simulation (The experimental setup for thermal data acquisition is shown in Fig. 4A, while the corresponding thermal imaging results are presented in Fig. 4B). The results are shown in Fig. 4C. When overlaying the temporal temperature variations with the simulation graphs, despite some discrepancies, the T63 values showed close agreement: for the case without LIG, the simulation yielded 118.1 seconds while the experimental measurement was 110.7 seconds; for the case with LIG, the simulation predicted 53.9 seconds compared to the experimental value of 58.3 seconds. These results demonstrate that LIG effectively functions as a photothermal conversion and facilitates heat transfer to the internal liquid.

### v4

@@ Abstract
Robotic systems’ mobility is fundamentally constrained by their power sources and wiring requirements. While electrical actuation systems have achieved autonomy through battery power and wireless control, pneumatic actuators remain tethered to air supply sources. Liquid-to-gas phase change actuators utilizing low-boiling-point liquids offer a potential solution, though they typically require substantial thermal input through heating elements that maintain electrical dependencies. External heat sources, particularly light energy, present an alternative for terrestrial applications. However, despite their optical transparency, silicone-based materials have a high volumetric heat capacity and low thermal conductivity, which limits efficient photothermal energy transfer. Previous attempts to address this issue through the incorporation of graphene or metallic powder have compromised material properties, including reduced transparency and altered elastic moduli. Inspired by the tapetum lucidum structure found in the eyes of nocturnal animals, which enables efficient light utilization in low-light conditions, this study proposes a novel anisotropic bilayer soft actuator incorporating Laser-Induced Graphene (LIG) on the inner surface of the light-irradiated silicone layer. This creates an anisotropic structure with enhanced photothermal conversion capabilities while maintaining the advantageous properties of silicone. Comparative analysis demonstrates that the proposed actuator exhibits significantly higher photo-induced bending efficiency than conventional silicone-based actuators. The response time improved by 54%, decreasing from 142 seconds for pure silicone to 65 seconds, with recovery response time showing a 48% improvement. This design maintains the silicone’s transparency and flexibility while utilizing LIG, which can be fabricated under ambient conditions, facilitating manufacturing and diverse applications.
@@ 1 Introduction
Currently, most robotic actuators rely on electrical energy. For example, although pneumatic actuators use air as the driving source, they rely on electrical energy to control and supply pressed air [1, 2]. Therefore, they cannot be considered completely independent of electrical energy.
For centuries, people have drawn upon a wide array of energy sources to enhance their lives. Exploring alternatives to electrical power for driving soft robots can greatly broaden the range of actuation techniques available in robotics.
Several methods have been proposed for using heating methods, such as laser irradiation [3, 4], microwave [5], heatsink [6] and electromagnetic [7, 8]. Kim et al. developed an electrothermal pneumatic actuator using nanofiber mats, achieving a lightweight, soft actuator that can operate without external devices by utilizing liquid evaporation [9]. However, this approach faces issues such as high power dependency and slow response times.
Various proposals have been made, starting with actuators like pouch motors [10, 11, 12, 13] that utilize liquid-to-gas phase change, including pneumatic actuators driven by ambient temperature [14]. However, these actuators operate relatively slowly, with many functioning at frequencies below 0.1Hz. This is due to significant losses in converting environmental energy into driving energy, heat transfer, and thermal conductivity issues. Sogabe et al.[15] developed a small soft actuator that operates in environments rich in thermal energy, such as warm water. This innovation improved the actuation time for bending to around one second using only environmental energy. However, this actuator leverages the thermal properties of the surrounding water (specific heat capacity of 4186 $\rm{J/kg\cdot K}$[16]) and is thus limited to underwater use.
For broader environmental applications, systems capable of operating on land are required. Thermal energy available in natural environments includes not only water temperature or body heat[17] but also ambient thermal gradients, which have been utilized for shape memory actuation and energy harvesting in soft robotic systems[18]. In addition to these, other strategies harness photothermal conversion using sunlight as an energy source.
A solar-driven approach utilizes a selective absorber film and a low-boiling-point fluid to generate internal pressure through sunlight-induced evaporation, enabling untethered soft robotic actuation [19]. Shao et al. proposed an actuator that absorbs environmental light energy through blackbody radiation by mixing graphite into silicone, converting it into thermal energy, and driving it using the liquid-to-gas phase change (methanol)[20].
However, using this approach leads to reduced actuator transparency because of the graphite content. When powders such as graphite are incorporated, they decrease Young’s modulus and impose limitations on both the actuator’s form and performance capabilities. For this reason, developing an efficient environmental energy absorption mechanism becomes essential - one that allows even highly transparent silicone actuators to function effectively. Our work presents bio-inspired mechanisms designed to efficiently capture environmental energy for actuator operation.
@@
Nocturnal animals rely on extremely low ambient light from the moon and stars to hunt and avoid predators. To maximize the capture of this scarce light, many nocturnal species possess a specialized reflective layer behind the retina called the tapetum lucidum (Fig. 1A). By reflecting photons that have passed through the photoreceptors back into the retina, the tapetum lucidum significantly enhances photoreceptor stimulation under dim conditions.
Conventional silicone-layer-only liquid-to-gas phase-change actuators absorb only a fraction of incident light, resulting in inefficient photothermal conversion. To improve light utilization efficiency, we use laser-induced graphene (LIG) to absorb light transmitted through the silicone layer and convert it into heat, analogous to how the tapetum lucidum recovers and reuses light that would otherwise be lost.
LIG has become an important area of research in direct laser writing techniques. In this process, organic materials like polyimide are converted into porous, three-dimensional graphene-like structures through rapid thermal carbonization (pyrolysis) using $\rm{CO_{2}}$ laser irradiation [21, 22]. The resulting porous network efficiently absorbs incident photons across a broad spectrum and converts them to thermal energy without requiring external wiring or electrical connections.
Due to its high electrical conductivity and patternability, LIG has attracted considerable interest in flexible electronics and multifunctional sensing devices [23, 24]. While most research has concentrated on electro-responsive or sensing applications, these studies reveal the expanding capabilities of LIG-based platforms. Wang et al. proposed light-driven robots utilizing the photothermal Marangoni effect [25]. Furthermore, Dallinger et al. transferred LIG onto PDMS-based hydrogel sheets and achieved rapid, humidity-driven bending under light irradiation [26]. These properties make LIG a promising candidate for bioelectronics and soft actuator applications [27].
Building on these advances, we extended LIG’s application to liquid-to-gas phase-change actuators. To enable efficient light-driven actuation, we added a LIG layer to conventional silicone soft actuators [15], creating an anisotropic bilayer structure in which the LIG "tapetum" lies on the light-incident surface beneath highly transparent silicone. This design captures any light that passes through the silicone and converts it to thermal energy, which is then transferred to the internal low-boiling-point fluid, causing vaporization and controlled bending of the silicone actuator.
In this study, we detail the fabrication process of the proposed actuator and evaluate its performance characteristics. Finally, to demonstrate practical applicability, we developed a multi-finger robot in which multiple LIG-enhanced actuators operate synchronously in response to light stimuli.
@@ 2.1 Concept
Figure 1 illustrates the conceptual design of our proposed actuator. Conventional silicone-layer-only liquid-to-gas phase-change actuators absorb only a fraction of incident light, leading to inefficient photothermal conversion. By analogy to the tapetum lucidum’s function of recovering and reusing otherwise lost light, our novel approach adopts a layered wall architecture to maximize light absorption. Specifically, we incorporate a Laser-Induced Graphene (LIG) layer on the light-incident surface of a highly transparent silicone elastomer (Fig. 1A). This LIG "tapetum" captures any light that passes through the silicone, converting it into thermal energy via photothermal conversion to efficiently vaporize and expand the internal low-boiling-point liquid.
The actuation mechanism of the proposed device is depicted in Fig. 1B. Under external light irradiation, the LIG layer heats up and transfers thermal energy to the internal working fluid (Opteon SF33, Chemours Company; $\rm{Cis\text{-}CF_{3}CH=CHCF_{3}}$), causing it to vaporize. Because the silicone walls are purposely asymmetric—thinner on the LIG side—the actuator bends directionally toward the LIG layer when the fluid expands. The operating temperature is 33C (above room temperature, 25C), selected to match the boiling point of the chosen fluorinated liquid.
Figure 1C shows a photograph of the fabricated actuator, which measures 20 mm in length with a flattened elliptical cross-section of approximately 5 mm by 4 mm. As illustrated in Fig. 1D, our bilayer design achieves efficient heat transfer without embedding additional conductive or photothermal fillers into the silicone walls.
The LIG layer is applied to only a small portion of the actuator’s interior, preserving transparency across most of the device’s area. This approach leaves most of the silicone structure clear while still taking advantage of LIG’s heat-generating properties when exposed to light. This design maintains the actuator’s lightweight and flexible nature while leveraging LIG’s photothermal properties and facilitating easy fabrication under ambient conditions.
@@ 2.2.1 LIG Transfer Process onto Silicone
The LIG patterns are generated on Kapton tape (3M Polyimide Film Tape 5413, 3M Company, Minnesota, USA) adhered to a slide glass. The polyimide base of the Kapton tape enables LIG formation through laser irradiation. We utilized a laser system (xTool F1, Makeblock Co., Ltd., Guangdong, China), which is equipped with a 455 nm blue diode laser (rated power: 10 W). For LIG fabrication, the laser was operated at 40% power with a scanning speed of 200 mm/s in unidirectional mode [28].
@@ 2.3 LIG-transferred soft actuator fabrication process
Figure 2 illustrates the fabrication of a soft actuator with the LIG layer next to the liquid interface, consisting of five steps (i–v). First, a LIG pattern (2.6 mm $\times$ 40 mm) is created with the specified laser parameters. Next step, as shown in Fig. 2(i), the mold and the printed LIG pattern are superimposed and fixed in place. Next, a metallic rod (2.5 mm diameter) is placed in the mold, and silicone is poured and cured (Fig. 2(ii)). In Fig. 2(iii), the tubular silicone structure with integrated LIG is removed. Since the LIG is on the exterior, the tube is inverted to position LIG inside (Fig. 2(iv)). In Fig. 2(v), uncured silicone seals one tube end, yielding an actuator with an open terminus. Finally, 40 µL of low-boiling-point liquid (Opteon SF33, Chemours, USA) is added to the cavity, and the end is sealed with a silicone-coated clip.
Additionally, the silicone sheets (10 mm square with transferred LIG patterns) were fabricated using the same process with molds adjusted to achieve a thickness of 1.0 mm. These molds were fabricated using a Formlabs Form 3 printer (Formlabs Inc., MA, USA). Clear Resin v4 (Formlabs Inc.) was used as the mold production resin material. We evaluated temperature response to determine whether placing silicone externally and LIG internally affects photothermal efficiency. A reflector lamp (RF100 V54 WD, 100 V, 54 W; Panasonic, Japan) served as the light source. Temperature changes on the silicone sheet were measured with a thermal camera (OPTPI64 ILTO15 T090, Optris GmbH,Berlin, Germany), and illuminance was measured using an illuminance meter (LX20, Sanwa Electric Instrument Co., Ltd., Tokyo, Japan).
Our findings revealed that positioning LIG on the interior accelerated the temperature rise at the liquid-contact interface (Supplementary Fig. 1). The initial temperature change rate during the first 5 seconds after light irradiation was 0.71 K/s when irradiated from the silicone side, compared to 0.66 K/s when irradiated from the LIG side (reference value: approximately 0.13 K/s for silicone-only sheet). This phenomenon can be attributed to silicone’s high volumetric heat capacity and low thermal conductivity, which impede the efficient transfer of heat generated by LIG to the liquid-contact interface.
@@ 2.4 Mechanical and chemical durability of LIG-transferred silicone
To assess the mechanical integrity and interfacial bonding of LIG transferred onto silicone elastomers, we exposed samples to operational stresses. The experimental protocol included uniaxial stretching, torsional deformation, and chemical exposure tests using Opteon SF33. Test specimens consisted of silicone elastomers with 10 mm $\times$ 10 mm LIG patterns. Samples underwent biaxial elongation to 200% strain through 100 loading cycles ( 3A). Twisting testing involved $\pm$ 180°rotations repeated over 100 cycles ( 3B). Chemical compatibility was examined through 5-minute immersion in Opteon SF33 ( 3C). Adhesion was quantified using the tape peel-off test[29, 30]. After conditioning, 3M Scotch Tape was applied, equilibrated, and then removed from the substrate surface. Pre- and post-treatment samples were analyzed by scanning electron microscopy (SEM) to identify delamination. Removed tape sections were examined LIG debris.
Figure 3 presents SEM images of LIG surfaces before and after treatment protocols. Imaging was performed using a field-emission SEM (JSM-7800F-PRIME, JEOL Ltd., Tokyo) at 5 kV accelerating voltage. Specimens received osmium coating (Tennant 20, Meiwafosis Co., Ltd., Tokyo) prior to examination. SEM analysis revealed negligible morphological alterations following stretching, torsional, or immersion treatments. The tape peel-off test produced no observable delamination or LIG structural modification, with no detectable transfer residue on tape surfaces. These findings demonstrate robust LIG–elastomer interfacial bonding sufficient for anticipated actuator operating conditions.
@@ 3 Heat transfer analysis
Here, $\sigma=5.67037442\times 10^{-8}~{}\mathrm{W\cdot m^{-2}\cdot K^{-4}}$ is the Stefan-Boltzmann constant; $\theta_{s}$ is the temperature of the silicone wall; $A$ is the heat transfer area; and $\epsilon_{h}$ and $\epsilon_{s}$ are the emissivities of the heat source and the silicone wall, respectively [31]. The thermal conductivity of the silicone, $\lambda_{s}$, and its thickness, $x$, are given in Table 1 [32].
@@
The parameters used in the simulation are summarized in Table 1[33, 34].
@@
For comparison with simulation results, we measured the temperature at the liquid contact surface using a thermal camera with silicone sheets of the same thickness as in the simulation (The experimental setup for thermal data acquisition is shown in Fig. 4A, while the corresponding thermal imaging results are presented in Fig. 4B). The results are shown in Fig. 4C. When overlaying the temporal temperature variations with the simulation graphs, despite some discrepancies, the 63% response time values showed close agreement: for the case without LIG, the simulation yielded 100.0 seconds while the experimental measurement was 103.9 seconds; for the case with LIG, the simulation predicted 54.9 seconds compared to the experimental value of 62.6 seconds. These results demonstrate that LIG effectively functions as a photothermal conversion and facilitates heat transfer to the internal liquid.
The LIG–silicone actuator cooled more rapidly because it reached a higher peak temperature (56.4C vs. 36.7 C), creating a larger $\delta$ T relative to the ambient ( 25C). According to Newton’s law of cooling, this greater temperature difference increases the rate of heat loss. Furthermore, the porous LIG structure increases both surface area and thermal conductivity, which also accelerates dissipation.